OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

A parallel line implies a slope that is equal to the given equation (in this case, 1). The general form of the equation here will be y = x + b. Sub in the point for x and y to find b. After finding b, you will plug that into your standard form to get y = x + b (where b is the number you just found).

OpenStudy (anonymous):

y = x - 8 is right answer because slope of given line is one and lines are paraller so slopes will be same so by using point slope form y-y1=m(x-x1) substitute value